OpenStudy (anonymous):

How does the slope of the left branch relate to the slope of the right branch? How does "a" relate to the slope of the branches? y= 2|x-3|- 4 y=-|x-3|-4 y=|x-3|-4 y= -2|x+3|-4

OpenStudy (anonymous):

i think both the slopes are equal...

OpenStudy (inkyvoyd):

The slopes are not equal.

OpenStudy (anonymous):

nope the left hand slope is -2 and the right hand slope is 2

OpenStudy (anonymous):

to solve this....you need to get rid of the absolute value first. So you need to address 2 circumstances. One when whats inside the absolute value is positive: so y=2(x-3) -4 y=2x-10 The next circumstance is when what is inside the absolute value is negative. so y=-2(x-3)-4 This gives you y=-2x+2